OpenStudy (anonymous):

A projectile is fired horizontally from a gun that is 45.0 m above flat ground, emerging from the gun with a speed of 250 m/s. (a) How long does the projectile remain in the air? (b) At what horizontal distance from the firing point does it strike the ground? (c) What is the magnitude of the vertical component of its velocity as it strikes the ground?

OpenStudy (phi):

Do you know how far an object falls due to gravity?

OpenStudy (anonymous):

9.8m/s^2

OpenStudy (phi):

That is acceleration. you need the formula for the distance an object falls when its acceleration is 9.8 m/s^2

OpenStudy (anonymous):

is it y-y0= voy(-1/2gt^2)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i used it but get the wrong answer

OpenStudy (phi):

distance = 1/2 g t^2 is how I think of it. The object starts at 45 m high and falls. so solve 45 = 1/2 * 9.8 m/s^2 * t^2

OpenStudy (anonymous):

oh so you don't include the 250m/s ahhhhh

OpenStudy (anonymous):

i guess that is x velocity

OpenStudy (anonymous):

so thats why...bah

OpenStudy (phi):

If you shot down (or up) , then you have to account for that, but horizontal the object goes sideways and falls at the same time.

OpenStudy (anonymous):

excellent. thanks

OpenStudy (anonymous):

ok do you know how to approach c?

OpenStudy (phi):

V= a*t or in this case V= gt

OpenStudy (phi):

For (a) you should get something very close to 3 sec (3 if g=10 exactly) (b) close to 750 m (c) close to 30 m/s